OpenStudy (anonymous):

Find the exact solutions of the given equation in the interval [0,2π]: sin^2x + sin2x = -cos^2x + 1

OpenStudy (jamesj):

well, first of all, remember that \[ \sin^2 x + \cos^2 x = 1\] Use that to simplify your problem!

OpenStudy (mertsj):

\[\sin ^2x+\sin 2x+\cos ^2x-1=0\] \[1+\sin 2x-1=0\] \[\sin 2x=0\]

OpenStudy (mertsj):

Take it from there.

OpenStudy (anonymous):

ok, i will give it a try

OpenStudy (anonymous):

i know that 0 would work for a solution, but what is the best way to find the rest of the choices?

OpenStudy (anonymous):